40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13413     Entry time: Tue Nov 7 22:56:21 2017     In reply to: 13412     Reply to this: 13414
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking recovered 

I hadn't re-locked the DRMI after the work on the AS55 demod board. Tonight, I was able to recover the DRMI locking with the old settings.

The feature in the PRCL spectrum (uncalibrated, y-axis is cts/rtHz) at ~1.6kHz is mysterious, I wonder what that's about.

Wasted some time tonight futzing around with various settings because I couldn't catch a DRMI lock, thinking I may have to re-tune demod phases etc given that I've been mucking around the LSC rack a fair bit. But fortunately, the problem turned out to be that the correct feedforward filters were not enabled in the angular feedforward path (seems like these are not SDF monitored). Clue was that there was more angular motion of the POP spot on the CCD than I'm used to seeing, even in the PRMI carrier lock.

After fixing this, lock was acquired within seconds, and the locks are as robust as I remember them - I just broke one after ~20mins locked because I went into the lab. I've been putting off looking at this angular feedforward stuff and trying out some ideas rana suggested, seems like it can be really useful.

As part of the pre-lock work, I dither aligned arms, and then ran the PRCL/MICH dithers as well, following which I re-centered the ITM, PRM and BS Oplevspots onto their respective QPDs - they have not been centered for a couple of months now.

I'm now going to try and measure some other couplings like PSL RIN->MICH, Marconi phase noise->MICH etc.

 

Attachment 1: DRMI_7Nov20178.png  48 kB  | Hide | Hide all
DRMI_7Nov20178.png
ELOG V3.1.3-